# Measuring the Measures: Discriminative Capacity of Representational Similarity Metrics Across Model Families

### v1

@@ 3 Results
Our systematic evaluation reveals clear differences in the discriminative capacity of representational similarity metrics (Figures 1, 2). RSA demonstrates the strongest overall separability with a d-prime of 3.79, closely followed by SoftMatch ($d^{\prime}=3.61$), then Procrustes ($d^{\prime}=3.20$), with Linear Predictivity showing substantially weaker discrimination ($d^{\prime}=2.38$). The pattern is particularly pronounced for silhouette coefficients, where RSA achieves a notably high score of 0.51—indicating robust within-family clustering—while the mapping-based metrics show progressively weaker clustering: SoftMatch (0.29), Procrustes (0.21), and Linear Predictivity (0.13). ROC-AUC analysis confirms this hierarchy (Figure 3: RSA (0.912) and SoftMatch (0.909) achieve superior classification of within- versus between-family pairs, followed by Procrustes (0.899) and Linear Predictivity (0.811).

### v2

@@ 3 Results
Our systematic evaluation reveals clear differences in the discriminative capacity of representational similarity metrics (Figures 1, 2). RSA demonstrates the strongest overall separability with a d-prime of 3.95, closely followed by SoftMatch ($d^{\prime}=3.59$), then Procrustes ($d^{\prime}=3.18$), with Linear Predictivity showing substantially weaker discrimination ($d^{\prime}=2.33$). The pattern is particularly pronounced for silhouette coefficients, where RSA achieves a notably high score of 0.56—indicating robust within-family clustering—while the mapping-based metrics show progressively weaker clustering: SoftMatch (0.29), Procrustes (0.21), and Linear Predictivity (0.13). ROC-AUC analysis confirms this hierarchy (Figure 3: RSA (0.9257) and SoftMatch (0.8994) achieve superior classification of within- versus between-family pairs, followed by Procrustes (0.8979) and Linear Predictivity (0.8137).